integral curve of dx/y^2+yz+x^2= dy/y^2-xz+x^2=dz/z(x+y)


1
Expert's answer
2021-08-30T16:47:26-0400

We first subtract the first expression from the second and equate to the third expression

"\frac{dy-dx}{y^{2}+yz+x^{2}-y^{2}+xz-x^{2}}=\frac{dz}{z(x+y)}"

"\frac{dx-dy}{z(y+x)}=\frac{dz}{z(x+y)}"

Simplify to get

d(x-y)=1dz

Integrate both sides

"\int d(x-y)=\int 1dz"

x-y+C=z

z=x-y+C1 ,this is the first integral curve

Second we add the third expression to the second and then equate with the first expression

"\frac{dx}{y^{2}+yz+x^{2}}=\frac{dy+dz}{y^{2}-xz+x^{2}+xz+yz}"

Simplify to get

"\frac{dx}{y^{2}+yz+x^{2}}=\frac{dy+dz}{y^{2}+yz+x^{2}}"

Further simply by multiplying both sides with y2+yz+x2 to get

dx=dy+dz=d(y+z)

Integrate both sides

"\int dx=\int d(y+z)"

x+C2=y+z

z=x-y+C2

Thus the equation only has one integral curve: z=x-y+C

Which can also be written, z-x+y=C